# Collaborative-Online-Learning-Enabled Distributionally Robust Motion Control for Multi-Robot Systems

### v1

@@ 4.3 Ambiguity set compression
Although the ambiguity set (36) can be directly employed to construct the DR-CVaR constraint (10), (41a) implies that the number of the basic ambiguity sets $m_{\ell,t,k}$ increases rapidly with $k$, thus posing a computational challenge for its practical application. To address this, we introduce a compression scheme that limits the total number of basic ambiguity sets to a user-defined maximum $M_{\ell,t,k}$ by aggregating similar components.
@@
$\hat{\widetilde{\gamma}}^{n}=\sum\nolimits_{j\in\mathbb{M}^{n}}\hat{\gamma}^{j},\quad\widetilde{\theta}=\theta$ (52a)
@@
$\sum_{j\in\mathbb{M}^{n}}\overline{\gamma}^{n}\left\{\left(\varepsilon^{j}+2\beta^{j}\right)\hat{\Sigma}^{j}+3\left(\hat{\mu}^{j}-\hat{\widetilde{\mu}}^{n}\right)\left(\hat{\mu}^{j}-\hat{\widetilde{\mu}}^{n}\right)^{T}\right\}$
@@ 5.1 Inter-robot separating hyperplane computation
We note that even for the two separated convex sets $\hat{y}_{i}(k|t)\oplus\mathcal{R}_{i}$ and $\hat{y}_{j}(k|t)\oplus\mathcal{R}_{j}$, the hyperplane defined by (54) and (55) sometimes cannot split them into two half-spaces. In this case, we directly use the separating hyperplane from the previous time step, i.e., $\mathcal{H}_{i,j,t,k}^{rob}=\mathcal{H}_{i,j,t-1,k+1}^{rob}$. In particular, for the case k=K1k=K-1, since $\mathcal{H}_{i,j,t-1,K}^{rob}$ does not exist, we compute $\mathcal{H}_{i,j,t,K-1}^{rob}$ as the separating hyperplane between $\hat{y}_{i}(K-1|t)\oplus\mathcal{R}_{i}$ and $\hat{y}_{j}(K-1|t)\oplus\mathcal{R}_{j}$. In summary, robots $i$ and $j$ are always separated into two half-spaces.
@@ 7       $\{N^{dat}_{i,\ell,t-1},S^{tr}_{i,\ell,t-1}\}\leftarrow\{N^{dat}_{j,\ell,t-1},S^{tr}_{j,\ell,t-1}\}$;
11    $\mathcal{H}_{i,j,t,k}^{rob}\leftarrow$ Inter-robot separating hyperplane calculation (Section 5.1) ;
@@ 21          $\widetilde{\mathbb{D}}_{\ell,t,k}\leftarrowAmbiguity$ set compression;
26    $\mathcal{H}_{i,\ell,t,k}^{obs}\leftarrow$ Robot-obstacle separating hyperplane calculation (Section 5.2) ;
@@
5.3 The algorithm and performance analysis
The COOL-DRMC algorithm for robot $i$ at time $t$ is delineated in Algorithm 2. The robot $i$ first communicates with other robots to execute COOL, and then calculates the separating hyperplanes $\mathcal{H}_{i,j,t,k}^{rob}$ and $\mathcal{H}_{i,\ell,t,k}^{obs}$, and finally solves the optimization problem (11). The first step of $\bm{u^{*}_{i}(t)}$ is utilized as control input $u^{*}_{i}(0|t)$. Additionally, we provide the performance analysis of COOL-DRMC. Specifically, we theoretically establish the probabilistic collision avoidance guarantee and the long-term tracking performance guarantee of COOL-DRMC. The probabilistic collision avoidance guarantee is as follows.
@@ 5.3 The algorithm and performance analysis
Let optimization problem (11) is feasible at time $t$, then the control trajectory $y_{i}(t+1)=y_{i}(1|t)$ is collision avoidance with at least $\prod_{\ell=1}^{L}\alpha_{\ell}(\alpha^{u})^{L}$, where $\alpha^{u}$ is the user-defined confidence level. Furthermore, let the optimization problem (11) is feasible at each time t=0,,T1t=0,...,T-1, then the control trajectory $y_{i}(1),...,y_{i}(T)$ is collision-free with at least $\prod_{\ell=1}^{L}(\alpha_{\ell})^{T}(\alpha^{u})^{TL}$.
@@
The system model $f$ is continuous, and the stage cost $l_{i}$ and the terminal cost $q_{i}$ are continuous at the origin and satisfy $l_{i}(r_{i}^{x}(t),r_{i}^{u}(t),t)=0$ and $q_{i}(r_{i}^{x}(t),t)=0$. In addition, $l_{i}(x_{i}(t),u_{i}(t),t)\geq g(\|x_{i}(t)-r_{i}^{x}(t)\|)$ for all feasible $x_{i}(t)$, $u_{i}(t)$, where $g$ is a $\mathcal{K}_{\infty}$-function.
@@
There exists a robust invariant set denoted $\mathcal{X}_{i}^{f}(t+K)\subseteq\mathcal{X}$ such that for all $x_{i}(K|t)\in\mathcal{X}_{i}^{f}(t+K)$, there exists a terminal control law $u_{i}^{f}(K|t)=\kappa_{i}^{f}(x_{i}(K|t),t+K)\in\mathcal{U}$ entailing that $x_{i}^{f}(K+1|t)=f(x_{i}(K|t),u_{i}^{f}(K|t))\in\mathcal{X}_{i}^{f}(t+K+1)$ and $q_{i}(x_{i}^{f}(K+1|t),t+K+1)-q_{i}(x_{i}(K|t),t+K)\leq-l_{i}(x_{i}(K|t),u_{i}^{f}(K|t),t+K)$. Moreover, for all $x_{i}(K|t)\in\mathcal{X}_{i}^{f}(t+K)$ and $\ell\in\mathbb{N}_{++}^{L}$, $(y_{i}(K|t)\oplus\mathcal{R}_{i})\cap(\mathbb{W}_{\ell,t+K}\oplus\mathcal{O}_{\ell})=\emptyset$. And for all $y_{j}(K|t)\in\mathcal{X}_{j}^{f}(t+K)$ and $i\neq j$, $(y_{i}(K|t)\oplus\mathcal{R}_{i})\cap(y_{j}(K|t)\oplus\mathcal{R}_{j})=\emptyset$.
Assumption 18 imposes the condition $q_{i}(x_{i}^{f}(K+1|t),t+K+1)-q_{i}(x_{i}(K|t),t+K)\leq-l_{i}(x_{i}(K|t),u_{i}^{f}(K|t),t+K)$, which essentially requires the terminal control policy to ensure a strict decrease in the cost-to-go function. While this assumption guarantees long-term tracking stability, it can be overly restrictive in motion planning tasks involving complex environments. This is because randomly moving obstacles may frequently intrude into the reference path, making long-term tracking infeasible in essence. Fortunately, this assumption is only necessary for ensuring the long-term tracking performance. If we are instead concerned with recursive feasibility, this condition can be safely relaxed. In particular, the terminal set $\mathcal{X}^{f}_{i}$ can be computed as a robust control invariant set. For robotic systems, such a robust terminal set can be defined by enforcing the velocity to be zero and the position to lie in the safe region, i.e., the complement of the union of the support sets of all obstacles. This safe region is nonempty in virtually all motion planning scenarios (at least in all case studies in Section 7). Moreover, for linear robotic systems, less conservative terminal sets can be computed using tools such as Multi-Parametric Toolbox (MPT) Herceg et al. (2013), which allow for the construction of the maximal robust control invariant set inside the safe region, thereby reducing conservatism compared to a zero-velocity constraint. Readers can refer to Section IV.B of Batkovic et al. (2022) for further explanations and practical examples regarding Assumption 18.
Assumption 19.
For all $\ell$, $t$, and $k$, $\mathbb{D}_{\ell,t+1,k-1}\subseteq\mathbb{D}_{\ell,t,k}$ is satisfied.
Assumption 19 essentially assumes that the uncertainty cannot increase as additional information becomes available. A similar assumption is frequently encountered in the literature on stochastic MPC and distributionally robust MPC Coppens and Patrinos (2021); Batkovic et al. (2022). With the above three assumptions, we provide the long-term tracking performance guarantee of COOL-DRMC as follows, which is proven in Appendix F.
@@
Suppose Assumptions 17, 18, and 19 hold and the optimization problem (11) is feasible at $t=0$, then the following long-term tracking performance guarantee is satisfied.
@@ Appendix B Proof of Lemma 9
For ease of exposition, we use $\mu,\mu_{\ell,t,k}^{j}$ and $\Sigma,\Sigma_{\ell,t,k}^{j}$ represent the means and covariances of $p$ and $p_{\ell,t,k}^{j}$, respectively. As a result, the two conditions in $\mathbb{D}$ (25) can be stated as follows.
$(\mu-\hat{\mu})^{T}(\hat{\Sigma})^{-1}(\mu-\hat{\mu})\leq\beta$ (99)
$\Sigma+(\mu-\hat{\mu})(\mu-\hat{\mu})^{T}\preceq\varepsilon\hat{\Sigma}$ (100)
Note that since $\hat{\Sigma}$ is positive definite, equation (99) has the following three equivalent forms.
$(\mu-\hat{\mu})^{T}(\hat{\Sigma})^{-1}(\mu-\hat{\mu})\leq\beta$ (101a)
$\Longleftrightarrow$ $\left(\begin{array}[]{cc}\hat{\Sigma}&(\mu-\hat{\mu})\\ (\mu-\hat{\mu})^{T}&\beta\end{array}\right)\succeq 0$ (101d)
$\Longleftrightarrow$ $(\mu-\hat{\mu})(\mu-\hat{\mu})^{T}\preceq\beta\hat{\Sigma}$ (101e)
Equation (100) has the following two equivalent forms.
$\Sigma+(\mu-\hat{\mu})(\mu-\hat{\mu})^{T}\preceq\varepsilon\hat{\Sigma}$ (102a)
$\Longleftrightarrow\Sigma\preceq(\varepsilon-\beta)\hat{\Sigma}$ (102b)
Clearly, the two conditions in $\mathbb{D}_{\ell,t,k}^{j}$ (36) have the corresponding equivalent forms above, which we will not elaborate further. We are now ready to begin the proof. Based on the relationship between $p_{\ell,t,k}^{j}$ and $p$ in (87), we can obtain
@@
Therefore $p_{\ell,t,k}^{j}$ satisfies the first condition in $\mathbb{D}_{\ell,t,k}^{j}$. For the second condition in $\mathbb{D}_{\ell,t,k}^{j}$, we verify its form corresponding to (102b) as follows.
$\Sigma_{\ell,t,k}^{j}=\sum\nolimits_{i=1}^{m_{\ell,t}}k_{i}^{j}\Sigma\preceq\sum\nolimits_{i=1}^{m_{\ell,t}}k_{i}^{j}(\varepsilon-\beta)\hat{\Sigma}$ (110)
$\preceq\max_{i=1...m_{\ell,t}}\{\varepsilon-\beta\}\sum\nolimits_{i=1}^{m_{\ell,t}}k_{i}^{j}\hat{\Sigma}=(\varepsilon_{\ell,t,k}^{j}-\beta_{\ell,t,k}^{j})\hat{\Sigma}_{\ell,t,k}^{j}$

### v2

@@ 4.3 Ambiguity set compression
Since the reformulation of the DR-CVaR constraint (11) only requires an ambiguity set over the obstacle position, the ambiguity set (37) can be directly employed to construct (11). However, as indicated by (42a), the number of basic ambiguity sets $m_{\ell,t,k}$ grows rapidly with $k$, which leads to a high computational burden in practice. To address this, we introduce a compression scheme that limits the total number of basic ambiguity sets to a user-defined maximum $M_{\ell,t,k}$ by aggregating similar components.
@@
$\hat{\widetilde{\gamma}}^{n}=\sum\nolimits_{j\in\mathbb{M}^{n}}\hat{\gamma}^{j},\qquad\widetilde{\theta}=\theta$ (53a)
@@
$\sum_{j\in\mathbb{M}^{n}}\overline{\gamma}^{n}\left\{\left(\varepsilon^{j}+3\beta^{j}\right)\hat{\Sigma}^{j}+3\left(\hat{\mu}^{j}-\hat{\widetilde{\mu}}^{n}\right)\left(\hat{\mu}^{j}-\hat{\widetilde{\mu}}^{n}\right)^{T}\right\}$
@@ 5.1 Inter-robot separating hyperplane computation
We note that even when the two convex sets $\hat{y}_{i}(k|t)\oplus\mathcal{R}_{i}$ and $\hat{y}_{j}(k|t)\oplus\mathcal{R}_{j}$ are disjoint, the separating hyperplane obtained via the specific construction in (55) and (56) may fail to separate them. In this case, we directly use the separating hyperplane from the previous time step, i.e., $\mathcal{H}_{i,j,t,k}^{rob}=\mathcal{H}_{i,j,t-1,k+1}^{rob}$. In particular, for the case k=K1k=K-1, since $\mathcal{H}_{i,j,t-1,K}^{rob}$ does not exist, we compute $\mathcal{H}_{i,j,t,K-1}^{rob}$ as the separating hyperplane between $\hat{y}_{i}(K-1|t)\oplus\mathcal{R}_{i}$ and $\hat{y}_{j}(K-1|t)\oplus\mathcal{R}_{j}$. In summary, robots $i$ and $j$ are always separated into two half-spaces.
@@ 5.2 Robot-obstacle separating hyperplane computation
5.3 COOL-DRMC with a novel update scheme
Based on the results in Sections 5.1 and 5.2, separating hyperplanes can be computed to enforce collision-avoidance constraints and ensure the safety of the multi-robot system. However, the affine constraints induced by these hyperplanes may render the trajectory optimization problem (12) infeasible. On the one hand, even when the committed convex regions are geometrically separated, the hyperplanes constructed by the adopted specific procedure may fail to separate them into non-intersecting half-spaces, causing previously committed trajectories to violate the resulting affine constraints. On the other hand, the ambiguity sets $\mathbb{D}_{\ell,t}$ are updated online, and the resulting changes in uncertainty may lead to newly computed affine constraints that are incompatible with existing committed trajectories. Therefore, a safe scheme for updating separating hyperplanes is required. Before introducing this scheme, we first define the terminal set $\mathcal{X}_{i}^{f}$ as follows.
Definition 16 (Terminal set).
The terminal set $\mathcal{X}_{i}^{f}$ of robot $i$ is defined as a robust positively invariant set over $\mathcal{X}$. Specifically, for any $x_{i}(t)\in\mathcal{X}_{i}^{f}$, there exists a terminal control law $u_{i}(t)=\kappa_{i}^{f}(x_{i}(t))\in\mathcal{U}$ entailing that $x_{i}(t+1)=f(x_{i}(t),u_{i}(t))\in\mathcal{X}_{i}^{f}$. Moreover, for all $j\neq i$, $\big(h(\mathcal{X}_{i}^{f})\oplus\mathcal{R}_{i}\big)\cap\big(h(\mathcal{X}_{j}^{f})\oplus\mathcal{R}_{j}\big)=\emptyset$. And for all $\ell\in\mathbb{N}_{+}^{L}$, $\big(h(\mathcal{X}_{i}^{f})\oplus\mathcal{R}_{i}\big)\cap\big(\mathbb{W}_{\ell,\varphi}\oplus\mathcal{O}_{\ell}\big)=\emptyset$.
For robotic systems, such a robust terminal set can be defined by enforcing the velocity to be zero and the position to lie in the safe region, i.e., the complement of the union of the support sets of all obstacles. This safe region is nonempty in virtually all motion planning scenarios (at least in all case studies in Section 7). Moreover, for linear robotic systems, less conservative terminal sets can be computed using tools such as Multi-Parametric Toolbox (MPT) Herceg et al. (2013), which allows for the construction of the maximal robust control invariant set inside the safe region, thereby reducing conservatism compared to a zero-velocity constraint. Readers can refer to Section IV.B of Batkovic et al. (2022) for further explanations and practical examples. Next, we present the safe update schemes for the inter-robot separating hyperplanes and the robot–obstacle separating hyperplanes, respectively.
For the inter-robot case, at time $t$ we consider the update of the separating hyperplane $\mathcal{H}_{i,j,t,k}^{rob}$ between robots $i$ and $j$ at the predicted time step $k|t$. Robot $i$ checks whether the committed occupied regions $\hat{y}_{i}(k|t)\oplus\mathcal{R}_{i}$ and $\hat{y}_{j}(k|t)\oplus\mathcal{R}_{j}$ lie strictly on opposite sides of the newly computed separating hyperplane. If so, the newly computed hyperplane is adopted. Otherwise, for $k\in\mathbb{N}_{++}^{K-1}$, the separating hyperplane from the previous time step is retained, i.e., $\mathcal{H}_{i,j,t,k}^{rob}=\mathcal{H}_{i,j,t-1,k+1}^{rob}$. For $k=K$, since $\mathcal{H}_{i,j,t-1,K+1}^{rob}$ does not exist, a separating hyperplane is constructed based on the terminal occupied regions $h(\mathcal{X}_{i}^{f})\oplus\mathcal{R}_{i}$ and $h(\mathcal{X}_{j}^{f})\oplus\mathcal{R}_{j}$, which is guaranteed to exist by Definition 16. Robot $j$ follows the same update rule, ensuring synchronous updates of the separating hyperplane. As a result, the inter-robot separating hyperplane remains feasible with respect to the committed trajectories.
For the robot-obstacle case, we consider the update of the separating hyperplane $\mathcal{H}_{i,\ell,t,k}^{obs}$ between robot $i$ and obstacle $\ell$ at the predicted time step $k|t$. For observable obstacles $\ell\in\mathbb{O}_{i,t}$, robot $i$ verifies whether the committed occupied region $\hat{y}_{i}(k|t)\oplus\mathcal{R}_{i}$ satisfies the affine constraint (12h) induced by the newly computed separating hyperplane. If the constraint is satisfied, the new hyperplane is adopted. Otherwise, for $k\in\mathbb{N}_{++}^{K-1}$, the hyperplane from the previous time step is retained, i.e., $\mathcal{H}_{i,\ell,t,k}^{obs}=\mathcal{H}_{i,\ell,t-1,k+1}^{obs}$. For $k=K$, since $\mathcal{H}_{i,\ell,t-1,K+1}^{obs}$ does not exist, a separating hyperplane is constructed based on the terminal occupied region $h(\mathcal{X}_{i}^{f})\oplus\mathcal{R}_{i}$ and the overall occupied region of obstacle $\ell$, $\mathbb{W}_{\ell,\varphi}\oplus\mathcal{O}_{\ell}$, whose existence is ensured by Definition 16. For unobservable obstacles $\ell\notin\mathbb{O}_{i,t}$, the same update rule as in the case where the newly computed separating hyperplane violates the committed trajectory constraints is applied. Under these safe update rules, the robot-obstacle separating hyperplanes remains feasible with respect to the committed trajectories at all times.
@@ 7       $\{N^{dat}_{i,\ell,t-1},S^{tr}_{i,\ell,t-1}\}\leftarrow\{N^{dat}_{j,\ell,t-1},S^{tr}_{j,\ell,t-1}\}$;
11    $\mathcal{H}_{i,j,t,k}^{rob}\leftarrow$ Inter-robot separating hyperplane calculation (Section 5.1) and check whether it should be updated (Section 5.3);
@@ 21          $\widetilde{\mathbb{D}}_{\ell,t,k}\leftarrowAmbiguity$ set compression;
26    $\mathcal{H}_{i,\ell,t,k}^{obs}\leftarrow$ Robot-obstacle separating hyperplane calculation (Section 5.2) and check whether it should be updated (Section 5.3);
@@
With the above safe update rules in place, the COOL-DRMC algorithm for robot $i$ at time $t$ is delineated in Algorithm 2. The robot $i$ first communicates with other robots to execute COOL, and then calculates the separating hyperplanes $\mathcal{H}_{i,j,t,k}^{rob}$ and $\mathcal{H}_{i,\ell,t,k}^{obs}$, and finally solves the optimization problem (12). The first step of $\bm{u^{*}_{i}(t)}$ is utilized as control input $u^{*}_{i}(0|t)$. Additionally, we provide the performance analysis of COOL-DRMC. Specifically, we theoretically establish the probabilistic collision avoidance guarantee and the long-term tracking performance guarantee of COOL-DRMC. The probabilistic collision avoidance guarantee is as follows.
@@
Let 12) be feasible at time $t$, then the control trajectory $y_{i}(t+1)=y_{i}(1|t)$ is collision avoidance with at least $\prod_{\ell=1}^{L}(\alpha_{\ell}\alpha^{u})$. Furthermore, let (12) be feasible at each time t=0,,T1t=0,...,T-1, then the control trajectory $y_{i}(1),...,y_{i}(T)$ is collision-free with at least $\prod_{\ell=1}^{L}(\alpha_{\ell}\alpha^{u})^{T}$.
@@
The system model $f$ is continuous, and the stage cost $l_{i}$ and the terminal cost $q_{i}$ are continuous at the origin and satisfy $l_{i}(r_{i}^{x}(t),r_{i}^{u}(t))=0$ and $q_{i}(r_{i}^{x}(t))=0$. In addition, $l_{i}(x_{i}(t),u_{i}(t))\geq g(\|x_{i}(t)-r_{i}^{x}(t)\|)$ for all feasible $x_{i}(t)$, $u_{i}(t)$, where $g$ is a $\mathcal{K}_{\infty}$-function.
@@
[…]
Assumption 19 essentially requires the terminal control policy to ensure a strict decrease in the cost-to-go function. While this assumption guarantees long-term tracking stability, it can be overly restrictive in motion planning tasks involving complex environments. This is because randomly moving obstacles may frequently intrude into the reference path, making long-term tracking infeasible in essence. Fortunately, this assumption is only necessary for ensuring the long-term tracking performance. If we are instead concerned with recursive feasibility, this condition can be safely relaxed. With the above two assumptions, we provide the long-term tracking performance guarantee of COOL-DRMC as follows, which is proven in Appendix F.
@@
Suppose Assumptions 18 and 19 hold and the optimization problem (12) is feasible at $t=0$, then the following long-term tracking performance guarantee is satisfied.
@@ Appendix B Proof of Lemma 9
For ease of exposition, we use $\mu_{\ell,t}^{i},\mu_{\ell,t,k}^{j}$ and $\Sigma_{\ell,t}^{i},\Sigma_{\ell,t,k}^{j}$ represent the means and covariances of $p$ and $p_{\ell,t,k}^{j}$, respectively. As a result, the two conditions in $\mathbb{D}_{\ell,t}^{i}$ (26) can be stated as follows.
$(\mu_{\ell,t}^{i}-\hat{\mu}_{\ell,t}^{i})^{T}(\hat{\Sigma}_{\ell,t}^{i})^{-1}(\mu_{\ell,t}^{i}-\hat{\mu}_{\ell,t}^{i})\leq\beta_{\ell,t}^{i}$ (100)
$\Sigma_{\ell,t}^{i}+(\mu_{\ell,t}^{i}-\hat{\mu}_{\ell,t}^{i})(\mu_{\ell,t}^{i}-\hat{\mu}_{\ell,t}^{i})^{T}\preceq\varepsilon_{\ell,t}^{i}\hat{\Sigma}_{\ell,t}^{i}$ (101)
Note that since $\hat{\Sigma}_{\ell,t}^{i}$ is positive definite, equation (100) has the following three equivalent forms.
$(\mu_{\ell,t}^{i}-\hat{\mu}_{\ell,t}^{i})^{T}(\hat{\Sigma}_{\ell,t}^{i})^{-1}(\mu_{\ell,t}^{i}-\hat{\mu}_{\ell,t}^{i})\leq\beta_{\ell,t}^{i}$ (102a)
$\Longleftrightarrow$ $\left(\begin{array}[]{cc}\hat{\Sigma}_{\ell,t}^{i}&(\mu_{\ell,t}^{i}-\hat{\mu}_{\ell,t}^{i})\\ (\mu_{\ell,t}^{i}-\hat{\mu}_{\ell,t}^{i})^{T}&\beta_{\ell,t}^{i}\end{array}\right)\succeq 0$ (102d)
$\Longleftrightarrow$ $(\mu_{\ell,t}^{i}-\hat{\mu}_{\ell,t}^{i})(\mu_{\ell,t}^{i}-\hat{\mu}_{\ell,t}^{i})^{T}\preceq\beta_{\ell,t}^{i}\hat{\Sigma}_{\ell,t}^{i}$ (102e)
Furthermore, based on (102e), the sufficient and necessary conditions for (101) can be derived as follows.
$\Sigma_{\ell,t}^{i}\preceq(\varepsilon_{\ell,t}^{i}-\beta_{\ell,t}^{i})\hat{\Sigma}_{\ell,t}^{i}$ (103a)
$\Longrightarrow$ $\Sigma_{\ell,t}^{i}+(\mu_{\ell,t}^{i}-\hat{\mu}_{\ell,t}^{i})(\mu_{\ell,t}^{i}-\hat{\mu}_{\ell,t}^{i})^{T}\preceq\varepsilon_{\ell,t}^{i}\hat{\Sigma}_{\ell,t}^{i}$ (103b)
$\Longrightarrow$ $\Sigma_{\ell,t}^{i}\preceq\varepsilon_{\ell,t}^{i}\hat{\Sigma}_{\ell,t}^{i}$ (103c)
It is evident that (103a) and (103c) are, respectively, the sufficient and necessary conditions for (103b). Clearly, the two conditions in $\mathbb{D}_{\ell,t,k}^{j}$ (37) have the corresponding forms above, which we will not elaborate further. We are now ready to begin the proof. Based on the relationship between $p_{\ell,t,k}^{j}$ and $p$ in (88), we can obtain
@@
Therefore $p_{\ell,t,k}^{j}$ satisfies the first condition in $\mathbb{D}_{\ell,t,k}^{j}$. For the second condition in $\mathbb{D}_{\ell,t,k}^{j}$, we verify the corresponding sufficient condition in (103a) as follows.
$\Sigma_{\ell,t,k}^{j}=\sum\nolimits_{i=1}^{m_{\ell,t}}k_{i}^{j}\Sigma\preceq\sum\nolimits_{i=1}^{m_{\ell,t}}k_{i}^{j}\varepsilon_{\ell,t}^{i}\hat{\Sigma}_{\ell,t}^{i}$ (111)
$\preceq$ $\max_{i=1...m_{\ell,t}}\varepsilon_{\ell,t}^{i}\sum\nolimits_{i=1}^{m_{\ell,t}}k_{i}^{j}\hat{\Sigma}_{\ell,t}^{i}=(\varepsilon_{\ell,t,k}^{j}-\beta_{\ell,t,k}^{j})\hat{\Sigma}_{\ell,t,k}^{j}$